OpenStudy (4815162342):

A solution is made by dissolving 1.8 moles of sodium chloride (NaCl) in 0.198 kilograms of water. If the molal boiling point constant for water (Kb) is 0.51 °C/m, what would be the boiling point of this solution? Show all the steps taken to solve this problem. [Hint: water boils at 100°C]

OpenStudy (4815162342):

@sweetburger @Matt6288

OpenStudy (matt6288):

Did you try to solve it?

OpenStudy (4815162342):

yes, but I dont know where to start.

OpenStudy (4815162342):

Can you tell me the formula and see if i do it right?

OpenStudy (matt6288):

I think theres a formula 1 sec

OpenStudy (4815162342):

how do you do that?

OpenStudy (4815162342):

How do you do that

OpenStudy (4815162342):

How do you solve?

OpenStudy (matt6288):

Oh I get it now!

OpenStudy (matt6288):

m is the molality m = 1.8/0.198 = 9.091

OpenStudy (4815162342):

are you replying?

OpenStudy (4815162342):

is that the answer?

OpenStudy (4815162342):

Is that the answer?

OpenStudy (matt6288):

so this is how it suppose to look (1.8) (9.091 m) (0.51 °C/m) = 8.345538 + 100 = The boiling point of the solution is 108.3°C

OpenStudy (matt6288):

Anyways thats what I think it is